OpenStudy (anonymous):

Okay so I need this WITH EXPERIMENTAL PROBABILITY!!! How would I set it up with a spinner?????? You take a three question multiple choice test. Each question has four choices. You don't know any of the answers. What is the expirimental probability that you will guess exactly two out of three correctly?

OpenStudy (anonymous):

please look at your last post because the answer given was wrong

OpenStudy (anonymous):

i gave you the correct one.

OpenStudy (anonymous):

i will type here if you like.

OpenStudy (anonymous):

I did. I wasn't there for the lesson, so how would I use this online spinner to do this? THat last one i spun 100 times and recorded, but that wont work for this

OpenStudy (anonymous):

ok i guess you take a spinner that has 4 equal parts, one is correct and the rest are not. so \[\frac{3}{4}\] of the area is one color and \[\frac{1}{4}\] is a different color. spin 3 times and record the number of times you get it "right"

OpenStudy (anonymous):

then repeat

OpenStudy (anonymous):

in any case the "theoretical " probability is not \[\frac{3}{64}\] it is \[\frac{9}{64}\]

OpenStudy (anonymous):

Thank you so much! Is this only for math, I am assuming> THis whole website?

OpenStudy (anonymous):

i think there is other stuff. good luck!

OpenStudy (anonymous):

:)